Common Products has just made its first issue of stock. It raised $1.4 million by selling 50,000 shares of stock to the public. These are the only shares outstanding. The par value of each share was $3. Complete the following table:

Common Stock (par value):

Additional paid-in capital:

Retained Earnings:

Net Common Equity: $2,000,000
Answer #1

common stock par value = 50000 * 3 = 150000

additional paid in capital = amount raised - common stock par value

= 1400000 - 150000

= 1250000

retained earnings = Net common equity - common stock value - additional paid in capital

= 2000000 - 150000 - 1250000

= 600000